A rectangular garage is 9ft longer than it is wide. If the area of the garage is 630ft.^2, find the length and the width.
Answer by checkley77(12844)

L=W+9
LW=630
(W+9)W=630
W^2+9W-630=0
(W+30)(W-21)=0
W-21=0
W=21 ANS FOR THE WIDTH.
L=21+9=30 ANS FOR THE LENGTH.
PROOF:
30*21=630
630=630